A fountain of water is located at the center of a circular pool. Not wishing 

to get his feet wet, a student walks around the pool and measures its 

circumference to be 16.5 m. Next, the student stands at the edge of the pool 

and uses a protractor to gauge the angle of elevation at the bottom of the 

fountain to be 55.0°. How high is the fountain? 


1
Expert's answer
2021-02-15T17:42:27-0500

The water passes these R=16.5 meters in


"t=\frac{R}{v_x}=\frac{R}{v\text{ cos}\theta}."


On the other hand, a half of this this time is required to reach the highest point from the bottom:


"\frac t2=\frac{v_y}{g}=\frac{v\text{ sin}\theta}{g},"

substitute the time from the first equation:


"\frac{R}{v\text{ cos}\theta}=\frac{2v\text{ sin}\theta}{g},\\\space\\
v^2=\frac{Rg}{\text{sin}(2\theta)}."


During a half of this time, on the other hand, the water raises up, or makes the first part of the trajectory from the bottom to the highest point:

"h=\frac{v_y^2}{2g}=\frac{(v\text{ sin}\theta)^2}{2g}= \frac{\big[\frac{Rg}{\text{sin}(2\theta)}\big]\text{sin}^2\theta}{2g}=\\\space\\
=\frac R4\text{ tan}\theta=5.89\text{ m}."